A pickup can travel 300 miles in the same time that a car going 10 miles per hour faster can travel 350 miles. How fast is the pickup traveling?
Answer by richwmiller(9135)

A pickup can travel 300 miles in the same time that a car going 10 miles per hour faster can travel 350 miles. How fast is the pickup traveling?
p*t=300
(p+10)*t=350
p=60
t=5
60*5=300
70*5=350